OpenStudy (anonymous):

Simplify 8 -4/3 A. -1/16 B. 1/16 C.16 D.-16

OpenStudy (anonymous):

8 - 4/3 = 6+2/3 = 20/3

OpenStudy (dumbcow):

do you mean 8^(-4/3) ? = 1/(8^4/3) = 1/(2^4) = 1/16

OpenStudy (anonymous):

Thank you<3